# An Automated Grey Literature Extraction Tool for Software Engineering

### v1

@@ Abstract.
copyright: noneccs: Software and its engineering Software reliabilityccs: Software and its engineering Software maintenance and evolutionccs: Software and its engineering Software design engineeringccs: Computing methodologies Natural language processing
@@ 3. Tool Architecture and Workflow
GLiSE’S workflow consists of three steps that transform a user-provided prompt into a filtered and ranked corpus of grey literature items as illustrated by figure 1. For each step, we present the inputs and outputs, and give a short description of the processing performed.
@@ 3.1. Step 1: Query extraction
Input: A free‑text prompt describing the topic of interest (search intent).
@@ 3.3. Step 3: Relevance classification
Description: For each retrieved item, we compute embeddings (vector encoding of text semantics) for all relevant textual fields (e.g., title, snippet, README, issue body) using either text-embedding-3-small or text-embedding-3-large, depending on the user’s indications. We also embed the search intent itself. The embedding of each field is then compared with the search-intent embedding using vector-space operations (e.g. cosine distance, element-wise absolute difference), producing feature values that characterize their semantic similarity.
@@ 4. Data and Model Construction
4.1. Dataset Construction
To conduct our exploratory study and train our models, we first needed to construct an appropriate dataset. We began by defining a set of search intents, which we then transformed into concrete search queries using an LLM, following the same process used within the tool. These generated queries were sent to the APIs of the sources, and the retrieved results were collected and associated to their originating search intent. Then, for each provider, we manually curated and labeled the retrieved items according to whether they were relevant to their associated search intent or not. This process produced a set suitable for training and evaluating of our filtering models.
The composition and size of the resulting dataset are presented in Table 1.
@@ 4.1. Dataset Construction
The dataset is made available in the GLiSE repository (GLiSE2025).
@@ 4.3.1. Filtering Models Inputs
We experimented with different ways of transforming embeddings into model inputs. For each relevant textual field associated with retrieved results, we computed its embedding and paired it with the embedding of the search intent. For each such pair, we derived multiple features which we concatenated into input vectors for our ML models. We explored the following inputs combinations :
  • Cosine Distances
  • Euclidean Distances
  • L1 Distances
  • Cosine Distances and Euclidean Distances combined
  • All three distances combined
  • Element-wise absolute difference vectors
  • Element-wise product vectors
  • All features combined
@@ 4.3.5. LLM baseline
We also evaluated an LLM-based baseline (gpt-4o-2024-08-06) at temperature 0, using the most probable token for binary relevance prediction. We chose not to adopt this approach as it is slower, more costly, and generally inferior according to our results.
@@ 4.3.6. Final selection
Table 2. Evaluation of relevance classification models by web source, embedding configuration, and classifier type. Reported metrics correspond to the best-performing setup.
Source Representation Balanced Accuracy Precision F1 Score Recall Best ML Model/Dimensions/Features
@@ 5.1. Methodology
To measure and compare participants’ performance across both tasks, we used the following objective measures:
  • Time-To-First-Relevant (TTFR);
  • Items-To-10 (IT@10): The number of inspected items to find 10 relevant ones;
  • Screening/Research time to reach 10 relevant;
  • Coverage across sources (distinct providers represented in the final set).
@@ 5.2. Results and Analysis
SUS (0 to 100) 81.0
Usefulness (1 to 7) 6.0
Intention to reuse (1 to 7) 6.0
@@ 6. Threats to Validity
Internal validity: To mitigate risks of data leakage and overfitting while training our classifiers, we performed deduplication and shuffled the data before splitting, and used per-source partitions for training and evaluation. We ensured reproducibility by fixing random seeds. In the usability study, we also randomized task order and provided a short practice task to participants.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793304isbn: 979-8-4007-2474-9/2026/04ccs: Software and its engineering Software reliabilityccs: Software and its engineering Software maintenance and evolutionccs: Software and its engineering Software design engineeringccs: Computing methodologies Natural language processing
@@ 3. Tool Architecture and Workflow
In the first step, the LLM takes a prompt to generate queries. In the second step, the search engine API uses these queries to collect results from web sources and generate a “queries extraction” file. In the third step, this file is used by the ML classification model to generate a grey literature corpus
GLiSE’S workflow consists of 3 steps that transform a user-provided prompt into a filtered and ranked corpus of grey literature items as illustrated by figure 1. For each step, we present the inputs and outputs, and give a short description of the processing performed.
@@ 3.1. Step 1: Query extraction
Input: A free‑text prompt describing the search intent.
@@ 3.3. Step 3: Relevance classification
Description: For each retrieved item, we compute embeddings (vector encoding of text semantics) for all relevant textual fields (e.g., title, snippet, README) using either text-embedding-3-small or text-embedding-3-large, depending on the user’s indications. We also embed the search intent itself. The embedding of each field is then compared with the search-intent embedding using vector-space operations (e.g. cosine distance, element-wise absolute difference), producing feature values that characterize their semantic similarity.
@@ 4. Data and Model Construction
4.1. Dataset Construction
To conduct our exploratory study and train our models, we first needed to construct an appropriate dataset. We began by defining a set of search intents, which we then transformed into concrete search queries using an LLM, following the same process used within the tool. These generated queries were sent to the APIs of the sources, and the retrieved results were collected and associated to their originating search intent. Then, for each provider, we manually curated and labeled the retrieved items according to whether they were relevant to their associated search intent or not. This process produced a set suitable for training and evaluating of our filtering models. The resulting dataset is is available in the GLiSE repository (GLiSE2025) and its composition and size are presented in Table 1.
@@ 4.3.1. Filtering Models Inputs
We experimented with different ways of transforming embeddings into model inputs. For each relevant textual field associated with retrieved results, we computed its embedding and paired it with the embedding of the search intent. For each such pair, we derived multiple features which we concatenated into input vectors for our ML models. We explored the following inputs combinations : (i) Cosine Distances, (ii) Euclidean Distances, (iii) L1 Distances, (iv) Cosine Distances and Euclidean Distances combined, (v) All three distances combined, (vi) Element-wise absolute difference vectors, (vii) Element-wise product vectors, and (viii) All features combined.
@@ 4.3.5. LLM baseline
We also evaluated an LLM-based baseline (gpt-4o-2024-08-06) at temperature 0, using the most probable token for binary relevance prediction. We chose not to adopt this approach as it is slower, more costly, and inferior according to our baseline.
@@ 4.3.6. Final selection
Table 2. Selection of relevance classification models by web source, embedding configuration, and classifier type. Reported metrics correspond to the best-performing setup.
Source Representation Balanced Accuracy Precision Recall F1 Score Best ML Model/Dimensions/Features
@@ 5.1. Methodology
To measure and compare participants’ performance across both tasks, we used the following objective measures: (i) Time-To-First-Relevant (TTFR); (ii) Items-To-10 (IT@10) (the number of inspected items to find 10 relevant ones); (iii) Screening/Research time to reach 10 relevant; and (iv) Coverage across sources (distinct providers represented in the final set).
@@ 5.2. Results and Analysis
SUS (0 to 100) N/A 81.0
Usefulness (1 to 7) N/A 6.0
Intention to reuse (1 to 7) N/A 6.0
@@ 6. Threats to Validity
Internal validity: To mitigate risks of data leakage and overfitting while training our classifiers, we performed deduplication and shuffled the data before splitting, and used per-source partitions for training and evaluation. We ensured reproducibility by fixing random seeds. In the usability study, we also randomized task order and provided a short practice task to participants. Additionally, the reported classifier performance was used for model selection and should not be interpreted as an unbiased evaluation.